how do I solve 5d-1=-11
Found 2 solutions by Alan3354, shosta:
Answer by Alan3354(69443)

how do I solve 5d-1=-11
=====================
Add 1
Divide by 5

Answer by shosta(4)

to solve for d, try to get the d by itself on one side of the equal sign.
you can do this by getting rid of the 1 and 5 term on the same side as the d.
to get rid of the -1 we will add 1 to both sides
5d-1= -11
+1 +1
5d= -10
now to get rid of the 5 we have to divide it away, and whatever we do to one side of the equal sign we have to do to the other side
5d/5 = -10/5
d= -2 * we got the -2 because 5 goes into 10 2 times